OpenStudy (anonymous):

light of wavelength 1288 A is emitted when the hydrogen atoms drop from 5th to 3rd orbit.find the wavelength of ohoton emitted when electrons falls from 3rd to 2nd orbit????? plzz tell me

OpenStudy (nincompoop):

do you know what formula to use?

OpenStudy (nincompoop):

\[E=h \nu = h \left[ R \left( \frac{ 1 }{ n _{low}^{2} }-\frac{ 1 }{ n _{high}^{2} } \right) \right]\] h = planck's constant = \[h = 6.626 \times 10^{-34}\]

OpenStudy (anonymous):

yess i know formulaa tell me wht is the nxt step after thisss

OpenStudy (anonymous):

yess i know formulaa tell me wht is the nxt step after thisss

OpenStudy (nincompoop):

obtain the frequency \[\nu\]

OpenStudy (nincompoop):

\[R = 3.29 \times 10^{15}\]

OpenStudy (nincompoop):

\[\nu = R(\frac{ 1 }{ n _{low}^{2} }-\frac{ 1 }{ n _{high}^{2} })\] \[=3.29 \times 10^{15} \left( \frac{ 1 }{ 2^2 } - \frac{ 1 }{ 3^2 } \right)\]

OpenStudy (nincompoop):

it is necessary to obtain the frequency first since we need it to obtain the wavelength recall the formula \[\lambda \nu = c\] where: ƛ wavelength ν frequency (sometimes f is used) c speed of light 2.9979 x 10^8 m/s the speed of light is constant so we don't need to calculate that, thus leaving us only the frequency to solve in order the obtain the wavelength